OpenStudy (anonymous):

Solve x2 − 5x = 24. x = −3, x = 8 x = 3, x = −8 x = 6, x = −4 x = −6, x = 4 @hartnn

OpenStudy (anonymous):

its supposed to be x^2-5x=24 I will fan and give medal btw

OpenStudy (jskhupmang):

plug them in the equation and what ever that gets you zero will be the answer

OpenStudy (anonymous):

subtract 24 on both sides

OpenStudy (anonymous):

x^2-5x-24=0

OpenStudy (templeguy):

x=8,−3

OpenStudy (templeguy):

X=8,-3

OpenStudy (templeguy):

So A

OpenStudy (anonymous):

Thanks so much guys algebra is tricky for me lol

OpenStudy (anonymous):



OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

(x-8)(x+3)=0 x=8 and x=-3